OpenStudy (anonymous):

a≡b mod n means that a=b+kn, where k is an integer. What does this actually mean mean??

OpenStudy (ash2326):

a≡b mod n This means a=b+kn This actually means that the difference between a and b is an integral multiple of n For Example: 16≡10 mod 2 here the difference between 16 and 10=>16-10=6 6 is an integral multiple of 2 here k=3 \[16=10+3 \times 2\]

OpenStudy (anonymous):

Thanks for clarifying